OpenStudy (anonymous):

if Given : Line BC is congruent to Line EF Lines AC Is congruent to Line EGO Line AC is broken up in Segments AB =10 and BC=3 Line EG is broken up in Segments EF and FG How do I Prove FG= 10 With reasons??

OpenStudy (phi):

Do you mean Lines AC Is congruent to Line EG (not EGO) ? EF= BC reason: Given (BC congruent to EF) BC= 3 reason: given (length of BC is 3) EF= 3 reason: things equal to the same thing are equal to each other AC= EF reason: Given AC= AB+BC reason: a whole is the sum of its parts AC = 10+3 = 13 reason: given (AB=10 and BC=3) EG= EF+FG reason: whole is the sum of its parts EG= 3 + FG reason: EF = 3 EG= AC reason: given (AG congruent to EG) EG= 13 reason: AC=13. the parts must add up to the whole 13= 3+FG reason: 13= EG and 3+FG=13 (things = to the same thing are equal to each other 10= FG reason: